OpenStudy (anonymous):

Factor: \[\frac{ 2p-7 }{ 9 }^{3}\div \frac{ 7-2p }{ 3 }^{4}\]

OpenStudy (whpalmer4):

Are you sure that isn't \[\frac{(2p-7)^3}{9} \div\frac{(7-2p)^4}{3}\]?

OpenStudy (anonymous):

it is...

OpenStudy (whpalmer4):

Okay, what you wrote is very different...

OpenStudy (whpalmer4):

Do you remember how to divide by a fraction?

OpenStudy (anonymous):

you have to flip it because it changes to multiplication and then factor

OpenStudy (whpalmer4):

Flip it and multiply, right. If you flip and multiply, do you see any factors you can cancel?

OpenStudy (anonymous):

729 and 81 have 9 in common

OpenStudy (whpalmer4):

Where are you getting 729 and 81 from? If we flip and multiply, we have \[\frac{(2p-7)^3}{9}*\frac{3}{(7-2p)^4}\]What cancellation can you do right off the bat?

OpenStudy (anonymous):

9^3 and 3^4

OpenStudy (whpalmer4):

Unless you miswrote the problem in another way, the exponent isn't applied to the 3 and the 9...

OpenStudy (whpalmer4):

You said it was \[\frac{(2p-7)^3}{9} \div \frac{(7-2p)^4}{3}\]Are you now saying it is really \[(\frac{(2p-7)}{9})^3 \div (\frac{(7-2p)}{3})^4\]

OpenStudy (anonymous):



OpenStudy (whpalmer4):

Why did you say the other version was right?

OpenStudy (anonymous):

it doesnt matter now..thats the problem...can we just solve it?

OpenStudy (whpalmer4):

Well, no matter, I guess. After we flip and multiply, we get \[( \frac{2p-7}{9} )^3 * ( \frac{3}{7-2p} )^4\]What do you see to cancel there?

OpenStudy (anonymous):

you cant cancel until you multiply all that stuff....

OpenStudy (whpalmer4):

Sure you can. what about the 9 and the 3 to various powers?

OpenStudy (whpalmer4):

What is \[\frac{3^4}{9^3}\]

OpenStudy (anonymous):

big numbers

OpenStudy (whpalmer4):

Isn't that just \[\frac{3*3*3*3}{9*9*9}\]?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

And that simplifies to...

OpenStudy (whpalmer4):

\[9^3 = (3^2)^3 = 3^6\]

OpenStudy (whpalmer4):

\[\frac{3^4}{3^6} = 3^{4-6} = 3^{-2} = \frac{1}{3^2} = \frac{1}{9}\]

OpenStudy (whpalmer4):

Or if you're more into the visual, each pair of 3's in the numerator cancels out a 9 in the denominator, leaving \(\dfrac{1}{9}\)

OpenStudy (anonymous):

what about the 2p-7 and 7-2p?